y=-3/4x-1; 12,5- us this same equation to find parallel and perpendicular?
steps in writing the equation of a line parallel to the given line and passing through the given points
what is the equation of a line perpendicular to the given line and passing through the given points

Answer by stanbon(75887)

y=-3/4x-1; (12, 5)- use this same equation to find parallel and perpendicular?
steps in writing the equation of a line parallel to the given line and passing through the given points
---
The given line has slope = -3/4
----
Form of the parallel line:: y = mx + b, where m = -3/4, x = 12, y = 5::
Solve for "b"::
5 = (-3/4)12 + b
b = 5 + 9 = 14
---
Ans: y = (-3/4)x + 14
-----------------------------
what is the equation of a line perpendicular to the given line and passing through the given points
slope of all perpendicular lines must be 4/3
Form:: y = mx + b where m = 4/3 ; x = 12 ; y = 5
---
Solve for "b"::
5 = (3/4)12 + b
b = 5 -9 = -4
---
Ans: y = (3/4)x -4
------------------
Cheers,
Stan H.
-------------